Find the third angle.
y degrees, (y-20)degrees, _______________

Answer by umail08(87)

The sum of the angles of a triangle is 180 degrees, so:
Let Angle A = y degrees, Angle B = (y-20) degrees, Angle C = unknown
Angle A + angle B + Angle C = 180 degrees
Y + Y - 20 + Angle C = 180
2y – 20 + Angle C = 180
Angle C = 180 + 20 – 2y
Angle C = 200 – 2y